# Wildfire Risk Prediction: A Survey of Recent Advances Using Deep Learning Techniques

### v4

@@ 2.2 Wildfire Ignition and Burring Risk
where $P(C_{i})$ and $P(C_{i}|I)$ represent the ratio of the area of ignition unit $C_{i}$ to the total study area and the ratio of the number of ignitions to the total study area, respectively, while $P_{i}$ denotes the ratio of the number of ignitions to the number of units over a one-year span. To mitigate the issue of class imbalance, smaller areas were merged into larger ones:
@@
It is important to note that the studies mentioned above utilized ground-based wildfire databases rather than the more commonly used MODIS or VIIRS fire/hotspot data in wildfire risk prediction research. This may be due to the relatively low spatiotemporal resolution of these data, making it difficult to detect ignition points with very low intensity and duration. Benali et al. (2016) compared the consistency of wildfire field records in Portugal, Greece, Alaska, California, and southeastern Australia with MODIS active fire products (MCD14ML). They proposed a clustering algorithm to calculate the ignition and end dates of a wildfire in the MCD14ML product and used the Nash–Sutcliffe Model Efficiency Index to estimate the consistency between satellite-estimated fire dates and ground records. The study found that satellite data could not estimate the start and end dates of most fires smaller than 500 ha. Similarly, Coskuner (2022) compared the performance of MODIS and S-NPP VIIRS active fire or hotspot detection products, i.e., MCD14ML V6 and VNP14IMG, in detecting wildfires across five different land cover types in Turkey and compared these detections with records from Turkey’s ground-based wildfire database. The authors found that both MODIS and VIIRS had very low detection capabilities for small fires (<1 ha) (<3.5%). They also cited studies by Fusco et al. (2019) and Ying et al. (2019), showing that MODIS fire products had high omission rates in the United States and China. Specifically, Fusco et al. (2019) reported detection rates for MODIS products in the United States ranging from 3.5% to 23.4%, with detection rates increasing with fire size. Similarly, Ying et al. (2019) found that in Yunnan Province, China, only 11.10% of 5,145 ground-recorded wildfire events were detected by MODIS fire products, with most omissions attributed to the small size of the fires. These analyses suggest that the commonly used MODIS and VIIRS fire and hotspot products are not suitable for estimating ignition locations and events, nor for corresponding ignition risk estimation.
@@
In contrast to wildfire ignition risk, wildfire burning risk is more broadly defined as the risk of a fire encountering a particular location, influenced by both ignition and spread. Therefore, most studies do not distinguish between ignition and spread when defining wildfire burning risk but consider both simultaneously. In general, there are two methods for defining wildfire burning risk: using historical burn data through explicit or implicit methods, or defining it based on simulation or spread models (Finney, 1998, 2002; Scheller et al., 2007; Miller et al., 2008; Tolhurst et al., 2008; Tymstra et al., 2010; Chen et al., 2024d). These methods have been discussed in the previous sections of this chapter, so they will not be repeated here. It is important to note that in choosing between these two methods, (Miller and Ager, 2012) suggests that for smaller fire extents (e.g., 1-10 hectares), ignition risk assessment can accurately reflect the likelihood of fire occurrence. However, for larger fire extents (e.g., 20-50 km), it is necessary to consider using burning risk to represent wildfire risk.
@@ 3 Research area distribution
This paper counts 138 recent wildfire risk prediction works in various regions around the world. Fig. 2 illustrates the geographic distribution and quantity of wildfire prediction research globally. The figure indicates that North America and Asia are the most concentrated regions for wildfire prediction studies. The United States has the highest number of studies, totaling 71, followed by China and Australia, each with 18 studies. Additionally, Canada, Iran, and India exhibit considerable research activity, with 11, 14, and 6 studies, respectively.
@@
In Europe, the research is primarily concentrated in Western and Southern Europe. Spain, France, Greece, and Portugal have a notable number of studies, with 8, 5, 7, and 9 studies, respectively. Other countries such as Italy, Sweden, Serbia, and Turkey also have a smaller number of studies. Research on the African continent is relatively sparse, with South Africa, Algeria, Morocco, and Senegal each having one study. Similarly, South America has a limited number of studies, with Brazil and Chile each contributing one study. In Oceania, aside from Australia, no other countries have been recorded as having studies.
Furthermore, there are several studies categorized as global, totaling 18, and studies focused on the Mediterranean basin, totaling 6. These global and regional studies highlight the complexity and diversity of wildfire prediction across different climate zones and ecosystems.
Given the geographic distribution and quantity of research, future studies could focus on expanding research regions, promoting cross-regional studies, and enhancing data sharing and model validation. Specifically, although North America and Asia have concentrated research efforts, other regions such as Africa, South America, and parts of Europe have significantly fewer studies. Future research should aim to strengthen wildfire prediction studies in these underrepresented regions to address the geographic distribution gaps and emphasize the climate justice issue (Jones et al., 2024). Additionally, global and Mediterranean basin studies suggest that cross-regional research is crucial for understanding wildfire behavior under different ecosystems and climatic conditions. Enhancing international collaboration to conduct more comprehensive cross-regional studies can provide more robust prediction models and methods. Finally, many current studies rely on region-specific datasets, limiting the generalizability and applicability of the models. Future efforts should focus on promoting data sharing and openness, establishing unified data platforms, and conducting model validation across different regions to improve the reliability and applicability of prediction models.
@@ 4 Data
Essentially, wildfires are uncontrolled processes involving the continuous combustion of organic matter in landscapes (Duff et al., 2017). The ignition and burning processes of wildfires are influenced by factors such as fuel conditions, meteorological and climatic conditions, human interventions, topographical and hydrological characteristics, and historical wildfire occurrences (Fuller, 1991). Therefore, in this section, the data used in wildfire risk prediction are categorized into these types, and their components, definitions, and estimation or proxy methods are introduced.
@@ 4.1 Fuel Conditions
Duff et al. (2017) suggests that wildfire is a function of the characteristics of the fuel that sustains it. Compared to other factors influencing the ignition and spread of wildfires, such as meteorological conditions, climate, topography, and hydrology, fuel is relatively straightforward to control (Chuvieco et al., 2012). Some studies even posit that fuel is the only element in the landscape that can be altered to influence fire behavior (Thompson et al., 2013; Fernandes and Botelho, 2003). These studies highlight the importance of assessing fuel conditions in wildfire risk evaluations. Fuel refers to any combustible material; in the context of wildfire studies, this includes the combustible organic matter contained in both live and dead vegetation (McLauchlan et al., 2020; Leite et al., 2024). The influence of fuel conditions on wildfire potential, behavior, duration, and emissions is typically described using four variables: Fuel Moisture Content (FMC), fuel load, fuel type (fuel model), and fuel continuity. These parameters directly impact the likelihood of fire occurrence, its intensity, and its rate of spread.
@@ 4.1.1 Fuel moisture content
FMC is defined as the ratio of the weight of water (WoW) to the dry weight (DW) of the sample (Eq. 7)(Mbow et al., 2004; Chuvieco et al., 1999). FMC can be categorized into two types: Live Fuel Moisture Content (LFMC) and Dead Fuel Moisture Content (DFMC). It regulates the combustibility of the fuel (Yebra et al., 2013b) and is a significant factor influencing the spread rate of wildfires and the emissions produced during a fire (Fernandes et al., 2008; May et al., 2019).
$FMC=\frac{WoW}{DW}\times 100\%$ (7)
@@
In FMC, LFMC is influenced by various plant ecophysiological characteristics, adaptations, and environmental factors affecting transpiration, such as soil moisture and ambient humidity (Castro et al., 2003). In contrast, dead fuel differs from live fuel in that it cannot actively regulate its moisture content through water uptake via roots or water loss through stomatal transpiration. Instead, its moisture content is determined by fuel size, meteorological and microclimate conditions (Nolan et al., 2016), and the biochemical composition of the fuel (Viney, 1991a). These factors are generally determined through experiments or empirical data (Viney, 1991b; Ager et al., 2010). Dead fuel absorbs or loses moisture in response to environmental conditions and eventually reaches equilibrium with the surrounding atmosphere, known as Equilibrium Moisture Content (EMC). The time required to reach EMC is primarily determined by the size of the fuel (Catchpole et al., 2001). Fuel size can be classified into fine fuels, such as small twigs and leaves (1-hour fine fuels), coarser branches and small trunks (10-hour and 100-hour coarse fuels), large trunks and coarse branches (1000-hour fuels), dead grasses and forbs, as well as surface litter and fallen leaves. In some cases, soil is also considered a type of dead fuel, and its moisture content is measured (Rothermel et al., 1986). The moisture content of dead fuel, which is more influenced by environmental and weather factors, plays a decisive role in fire ignition and rapid spread at the beginning of the fire season, whereas the moisture content of live fuel has a greater impact on the persistence and extent of fire spread.
a) Dead Fuel Moisture Content
DFMC is a crucial input in nearly all wildfire risk assessment models (Matthews, 2013). For example, estimates of DFMC are integral to current operational fire prediction systems, including the Fire Weather Index (FWI) and the National Fire Danger Rating System (NFDRS) in the United States (Vitolo et al., 2020). Estimating DFMC involves calculating EMC based on factors like temperature, relative humidity, and fuel response time, combined with precipitation data and drying algorithms (Nieto et al., 2010). DFMC estimation methods can generally be divided into four categories: measurements at meteorological stations, field sampling, estimates based on meteorological data, and indirect retrieval using multi-source remote sensing data.
DFMC can be measured using standard wooden dowels installed 10-12 inches above the ground at meteorological stations (Rakhmatulina et al., 2021). However, due to the sparse network of DFMC measurement stations and the limited accuracy, it is challenging to cover large areas even with interpolation of atmospheric parameters. Similarly, field sampling of DFMC is highly accurate and provides data and theoretical support for physical process models and empirical modeling. However, field measurements of fuel properties like DFMC require consideration of terrain, vegetation type, ecosystem structure, climate type, and soil type, making it time-consuming, costly, and limited in scope (Duff et al., 2017). Therefore, studies combining field measurement data with station or numerical weather prediction and meteorological reanalysis data to develop empirical and mechanistic models for estimating FMC are more common (Matthews, 2013). Physical process models primarily consider energy and water balance equations, with the Nelson Dead Fuel Moisture Model, proposed by Nelson Jr (2000), being the most widely used for DFMC estimation. This model predicts DFMC by calculating heat conduction and moisture transfer between a wooden cylinder, the surrounding atmosphere, and external conditions, using hourly input data on temperature, humidity, radiation, and precipitation.
Additionally, McNorton and Di Giuseppe (2024) designed a new method for estimating DFMC by grouping fuels and simplifying the Nelson model. Other models used for DFMC estimation include bulk litter layer models, Byram’s diffusion equation-based models, and complete process-based models (Matthews, 2013). Empirical methods for estimating DFMC rely on observed relationships between measured DFMC and factors such as air temperature, relative humidity, wind speed, soil moisture content, and vapor pressure deficit. These data-driven methods include statistical techniques (Ferguson et al., 2002; Ray et al., 2010; Cawson et al., 2020; Rakhmatulina et al., 2021), but their primary limitation is that they are entirely data-driven, making it difficult to explain the relationship between driving factors like temperature, heat and water vapor fluxes, radiation, and precipitation, and DFMC. Consequently, some studies have started to combine data-driven approaches with physical process models (Fan and He, 2021; Fan et al., 2024).
Apart from the meteorological data-based FMC estimation methods mentioned above, indices such as the Cellulose Absorption Index (CAI), the Lignocellulose Absorption Index (LCAI), the Normalized Difference Tillage Index (NDTI), the Normalized Difference Vegetation Index (NDVI), and the Shortwave Infrared Normalized Difference Residue Index (SINDRI) have been shown to correlate with DFMC (ZORMPAS et al., 2017; Zacharakis and Tsihrintzis, 2023b). This forms the theoretical basis for DFMC estimation using remote sensing. However, research on DFMC estimation using multi-source remote sensing data is still limited (McCandless et al., 2020).
b) Live Fuel Moisture Content
In early research, LFMC was measured analogously to DFMC through field sampling and weighing, with broader area estimates obtained via interpolation or inversion techniques (Bianchi and Defossé, 2015). While locally precise, this method is neither scalable nor efficient for landscape or global applications due to its time-consuming and labor-intensive nature (Yebra et al., 2013a). Additionally, meteorological parameter-based LFMC estimation benefits from near real-time predictions but faces challenges due to the active regulation of water content by plants and their phenological differences (Pellizzaro et al., 2007; Ruffault et al., 2018).
As such, many studies favor multi-source geospatial data for large-scale and near real-time LFMC estimation, utilizing the growing capabilities of Earth observation technologies for high-resolution data (Rao et al., 2020a; Capps et al., 2021). These contemporary approaches utilize environmental variables such as precipitation and soil moisture, along with spectral, thermal infrared, and microwave data, often in combination with machine learning algorithms, to enhance prediction accuracy (Bowyer and Danson, 2004; Yebra et al., 2013a).
LFMC inversion methods based on multispectral reflectance data exploit water absorption bands in the near-infrared and shortwave infrared spectral regions. Drought stress leads to changes in chlorophyll concentration and leaf structure, enabling LFMC estimation through spectral reflectance relationships (Knipling, 1970; Yebra et al., 2013a). Physical models simulate these relationships, but they are complex; empirical methods using spectral indices like NDVI, EVI, and NDWI provide simpler alternatives and perform well at local scales (Hunt Jr et al., 1987; Fensholt and Sandholt, 2003). Studies have demonstrated the effectiveness of shortwave infrared bands and vegetation indices for LFMC estimation, though improvements through integrating meteorological data have also been suggested (Chuvieco et al., 2002; Myoung et al., 2018).
Thermal infrared bands, such as land surface temperature (LST), are also incorporated to enhance LFMC estimation since canopy temperature correlates with plant water stress (Chuvieco et al., 2004). Machine learning methods have further combined spectral and thermal data to model LFMC, revealing varying importance levels among predictors (McCandless et al., 2020).
Remote sensing methods often face limitations in lighting conditions and cloud coverage, prompting the use of microwave sensors for LFMC estimation due to their ability to penetrate clouds and sensitivity to water content (Rao et al., 2020b). Both active and passive microwave sensors have been employed, with studies showing promising results for vegetation monitoring and LFMC estimation (Dorigo et al., 2017; Lu and Wei, 2021). For example, integrating SAR and multispectral data with vegetation indices improves LFMC predictions, as shown in studies using Sentinel-1 and Landsat-8 data (Jia et al., 2019; Rao et al., 2020b).
Overall, advancements in combining multi-source data and modeling techniques present significant opportunities for accurate LFMC estimation. However, it is important to note that due to the complexity of real-world scenarios, LFMC estimation still faces significant challenges. These challenges include separating the effects of vegetation dry matter, soil moisture, and atmospheric and canopy structure interference on backscatter and optical reflectance and emissivity. Moreover, LFMC ground truth data remain scarce and sparsely distributed, hindering the development of data-driven algorithms.

### v5

@@ 2.2 Wildfire Ignition and Burning Risk
where $P(C_{i})$ and $P(C_{i}|I)$ represent the ratio of the area of ignition unit $C_{i}$ to the total study area and the ratio of the number of ignitions to the total study area, respectively, while $P(I|C_{i})$ denotes the ratio of the number of ignitions to the number of units over a one-year span. To mitigate the issue of class imbalance, smaller areas were merged into larger ones:
@@
It is important to note that the studies mentioned above utilized ground-based wildfire databases rather than the more commonly used MODIS or VIIRS fire/hotspot data in wildfire risk prediction research. This may be due to the relatively low spatiotemporal resolution of these data, making it difficult to detect ignition points with very low intensity and duration. Benali et al. (2016); Barber et al. (2024) compared the consistency of wildfire field records in Portugal, Greece, Alaska, California, and southeastern Australia with MODIS active fire products (MCD14ML). They proposed a clustering algorithm to calculate the ignition and end dates of a wildfire in the MCD14ML product and used the Nash–Sutcliffe Model Efficiency Index to estimate the consistency between satellite-estimated fire dates and ground records. The study found that satellite data could not estimate the start and end dates of most fires smaller than 500 ha. Similarly, Coskuner (2022) compared the performance of MODIS and S-NPP VIIRS active fire or hotspot detection products—specifically, MCD14ML V6 and VNP14IMG—in detecting wildfires across five different land cover types in Turkey. These detections were compared with records from Turkey’s ground-based wildfire database. The authors found that both MODIS and VIIRS had very low detection capabilities for small fires (<1 ha) (<3.5%). They also cited studies by Fusco et al. (2019) and Ying et al. (2019) showing that MODIS fire products had high omission rates in the United States and China. Specifically, Fusco et al. (2019) reported detection rates for MODIS products in the United States ranging from 3.5% to 23.4%, with detection rates increasing with fire size. Similarly, Ying et al. (2019) found that in Yunnan Province, China, only 11.10% of 5,145 ground-recorded wildfire events were detected by MODIS fire products, with most omissions attributed to the small size of the fires. These analyses suggest that the commonly used MODIS and VIIRS fire and hotspot products are not suitable for estimating ignition locations and events, nor for corresponding ignition risk estimation.
@@
In contrast to wildfire ignition risk, wildfire burning risk is more broadly defined as the risk of a fire encountering a particular location, influenced by both ignition and spread. Therefore, most studies do not distinguish between ignition and spread when defining wildfire burning risk but consider both simultaneously. In general, there are two methods for defining wildfire burning risk: using historical burn data through explicit or implicit methods, or defining it based on simulation or spread models (Finney, 1998, 2002; Scheller et al., 2007; Miller et al., 2008; Tolhurst et al., 2008; Tymstra et al., 2010; Chen et al., 2024c). These methods have been discussed in the previous sections of this chapter, so they will not be repeated here. It is important to note that in choosing between these two methods, Miller and Ager (2012) suggest that for smaller fire extents (e.g., 1-10 hectares), ignition risk assessment can accurately reflect the likelihood of fire occurrence. However, for larger fire extents (e.g., 20-50 $km^{2}$), it is necessary to consider using burning risk to represent wildfire risk.
@@ 3 Research Area distribution
This paper reviews 138 recent wildfire risk prediction studies conducted globally using data-driven methods. Figure 2 illustrates the geographic distribution and quantity of these studies. Among these, 18 studies address global-scale wildfire risks, while 6 focus specifically on the Mediterranean basin.
@@
At the continental scale, North America and Asia exhibit the highest levels of research activity. In North America, the United States accounts for 71 studies, followed by Canada with 11. In Asia, China leads with 18 studies, followed by Iran and India with 14 and 6 studies, respectively. In Europe, wildfire research is concentrated in the western and southern regions, with notable contributions from Spain, Portugal, Greece, and France, while smaller-scale studies have been conducted in Italy, Sweden, Serbia, and Turkey. In contrast, Africa has seen limited wildfire research, with only one study each from South Africa, Algeria, Morocco, and Senegal. Similarly, South America contributes minimally, with one study each from Brazil and Chile. In Oceania, research is dominated by Australia, which accounts for 18 studies, while no other countries in the region are represented.
The geographic concentration of wildfire research raises critical questions about whether these efforts align with global wildfire trends, particularly in terms of frequency, burned area, and carbon emissions. Africa, often referred to as the "continent of fire," has accounted for 70% of the global burned area over the past two decades, contributing approximately 29% of global carbon emissions (Giglio et al., 2013, 2018; de Oliveira-Júnior et al., 2024). Despite its prominence, Africa remains underrepresented in wildfire risk prediction studies, as shown in Figure  3. Similarly, South America, which contains approximately 864 million hectares of forest—accounting for 21% of the global forest area and 57% of the world’s primary forests (Food and Agriculture Organization of the United Nations (2011), FAO; Keenan et al., 2015)—has experienced a rising trend in wildfire frequency and burned area. This increase is primarily driven by deforestation and climate change. During the 2023–2024 wildfire season, the Amazon region, encompassing Brazil, Bolivia, Peru, and Colombia, experienced historically high burned areas, particularly in Amazonas, Loreto, La Paz, and Beni. Fire activity in June and October 2023 was the highest recorded since 1998. Southern South America, including Chile, also recorded significant wildfire activity during the same season. The burned area exceeded 4,000 square kilometers, marking the second largest in 20 years (Jones et al., 2024a). Despite these critical developments, data-driven wildfire research in South America remains sparse, as shown in Figure  3.
In northern boreal regions, Canada has been a focal point of wildfire research. However, studies on wildfire risk in Russia remain scarce, despite estimates suggesting that Siberian wildfire emissions surpass those of Canada (de Groot et al., 2013). Between 1998 and 2010, Russia’s average annual burned area and carbon emissions were 8.23 million hectares and 121 teragrams of carbon per year, respectively (Shvidenko et al., 2011). Recent studies indicate a growing prevalence of high-intensity, late-season fires in Siberian boreal forests, including the Arctic, with projections suggesting annual fire emissions may reach 325 teragrams of carbon by 2027 (Ponomarev et al., 2023). These findings underscore the urgent need to advance wildfire risk prediction in Russia’s northern regions.
In conclusion, the geographic distribution of wildfire research highlights significant disparities. While North America and Asia dominate wildfire studies, critical regions such as Africa, South America, and parts of Russia remain underrepresented. Addressing these gaps is vital to reducing geographic inequalities and advancing climate justice (Jones et al., 2024b). Expanding research to underrepresented regions, fostering cross-regional studies, and strengthening international collaboration are crucial to understanding wildfire behavior across diverse ecosystems and climates. Additionally, the reliance on localized datasets limits the generalizability of existing predictive models. Future efforts should prioritize the development of unified data platforms, facilitate open data sharing, and rigorously validate models across diverse regions to enhance their reliability and applicability.
@@ 4 Data
Essentially, wildfires are uncontrolled processes involving the continuous combustion of organic matter in landscapes (Duff et al., 2017). The ignition and burning processes of wildfires are influenced by factors such as fuel conditions, meteorological and climatic conditions, human interventions, topographical and hydrological characteristics, and historical wildfire occurrences (Fuller, 1991). Therefore, in this section, the data used in wildfire risk prediction are categorized into these types, and their components, definitions, and proxy methods are introduced.
It is important to note that most existing wildfire risk prediction studies rely on proxies, which can only indirectly reflect the state of driving factors. Therefore, in Section  8, we provide an overview of recent advancements in the calculation methods for these driving factors, with a particular focus on fuel conditions.
@@ 4.1 Fuel Conditions
Duff et al. (2017) suggests that wildfire behavior is largely dependent on the characteristics of the fuel that sustains it. Unlike other factors influencing wildfire ignition and spread, such as meteorological conditions, climate, topography, and hydrology, fuel can be more easily controlled (Chuvieco et al., 2012). Some studies even propose that fuel is the only landscape element that can be managed to directly affect fire behavior (Thompson et al., 2013; Fernandes and Botelho, 2003), emphasizing the importance of fuel conditions in wildfire risk evaluations. Fuel refers to any combustible material, including live and dead vegetation (McLauchlan et al., 2020; Leite et al., 2024), with four main variables affecting wildfire: Fuel Moisture Content (FMC), fuel load, fuel type, and fuel continuity. These parameters directly influence fire ignition, intensity, and spread rate.
@@ 4.1.1 Fuel Moisture Content
FMC refers to the ratio of the weight of water to the dry weight of a sample, as shown in Eq. 8 (Mbow et al., 2004; Chuvieco et al., 1999). FMC can be divided into Live Fuel Moisture Content (LFMC) and Dead Fuel Moisture Content (DFMC), both critical for determining fuel combustibility (Yebra et al., 2013b).
$FMC=\frac{\text{Weight of water in sample}}{\text{Dry weight of the sample}}\times 100\%$ (8)
@@
LFMC is influenced by plant ecophysiological factors and environmental conditions such as soil moisture and humidity, which regulate water uptake and transpiration (Castro et al., 2003). DFMC, on the other hand, is determined by fuel size, weather, and microclimate (Nolan et al., 2016), as well as the fuel’s biochemical composition (Viney, 1991). Dead fuel, unlike live fuel, cannot regulate moisture actively and reaches an equilibrium with the surrounding atmosphere known as Equilibrium Moisture Content (EMC). The time needed to reach EMC depends on the fuel size, which is classified into fine fuels like twigs and leaves (1-hour fuels), coarser branches (10-hour and 100-hour fuels), large trunks (1000-hour fuels), and surface litter (Catchpole et al., 2001). Soil moisture is also considered in some cases (Rothermel et al., 1986). DFMC plays a crucial role in fire ignition and rapid spread early in the fire season, while LFMC significantly impacts the persistence and spread of fires.
It is important to note that estimating FMC remains a significant challenge due to the complexity of real-world conditions. For example, accurately isolating the effects of vegetation dry matter, soil moisture, and atmospheric and canopy structural interference on backscatter, optical reflectance, and emissivity remains difficult. Additionally, the availability of in situ FMC measurements is still limited, with sparse spatial distribution, which constrains the development of data-driven algorithms.
To address these challenges, practical wildfire risk prediction systems often employ proxy parameters to represent DFMC and LFMC. For DFMC, these proxies include drought indices such as the Keetch-Byram Drought Index, meteorological variables like temperature and relative humidity, and composite indices such as the Fire Weather Index (FWI). These alternatives are easier to obtain and integrate into predictive models, while still capturing the indirect effects of DFMC on fire behavior (Vitolo et al., 2020; Rothermel et al., 1986).
For LFMC, proxy parameters such as vegetation indices (e.g., NDVI and its variants), meteorological variables (e.g., fuel moisture code, meteorological drought indices, and temperature), and composite indices (e.g., the Fire Weather Index) are commonly used, as shown in Figure 3. These proxies provide easier-to-obtain and more consistent inputs while indirectly capturing LFMC’s influence on wildfire behavior, thus offering a practical alternative for large-scale, long-term modeling efforts (Chuvieco et al., 2002; Dasgupta et al., 2007; Ruffault et al., 2018; Vinodkumar et al., 2021a; Krueger et al., 2022; Balaguer-Romano et al., 2022; Guk et al., 2024).